# Split Matching for Inductive Zero-shot Semantic Segmentation

### v2

@@ 4.1 Comparison with State-of-the-art methods
Table 1 presents an ablation study to evaluate the effectiveness of each proposed module: Split Matching (SM), Multi-scale Feature Enhancement (MFE), and the inclusion of additional randomly initialized queries. Starting from the baseline, the introduction of SM leads to a substantial improvement, increasing hIoU from 24.6 to 33.3 and uIoU from 20.0 to 30.8. Adding MFE further enhances the performance across all metrics, especially unseen classes (uIoU: 35.8). Finally, incorporating extra queries results in the best overall performance, achieving the highest scores in hIoU (36.6), sIoU (36.8), and uIoU (36.4). Note that our method achieves better performance even using a weaker backbone (ResNet101 vs. ViT-B).
@@ 4.2 Ablation Studies
To evaluate the effectiveness of our method, we conduct ablation studies on COCO-Stuff for 40K iterations using ResNet-50 as the backbone, with all hyperparameters unchanged. Additional experiments are provided in the Supplementary Materials due to space limitations.
@@
Ablations on Unseen Class Embedding. Table 4 presents an ablation study on the design of unseen class embeddings $\mathbf{C}_{u}$. Using raw dense features $\mathbf{F}_{u}$ from the backbone provides a reasonable baseline (uIoU: 32.1), while adding an MLP projection degrades performance, likely due to semantic distortion. In contrast, using CLS tokens from pseudo-masked regions yields the best results (uIoU: 36.4), suggesting that region-level CLS tokens better preserve semantic alignment. This also enables direct concatenation with seen-class text embeddings, maintaining cross-modal consistency without modality mismatch.
Ablations on structure of MFE. Table 4 presents the ablation study on the structure of the MFE module by comparing different normalization strategies: No Normalization (No Norm), Batch Normalization (BN), and Group Normalization (GN). Among the three, GN achieves the best performance across all metrics, yielding the highest hIoU (36.6), sIoU (36.8), and uIoU (36.4). These results demonstrate that incorporating Group Normalization into MFE significantly improves both seen and unseen segmentation performance.
@@ 4.3 Qualitative Analysis
Due to the space limitations, more results are in Supplementary Materials.
Visualization of candidate queries. Fig. 6 visualizes the predictions of candidate queries. Notably, these queries successfully activate on previously unannotated regions, enabling the model to localize unseen classes such as carrot and cow. This demonstrates that candidate queries can effectively discover latent categories and assign semantically correct class labels, even without explicit supervision.
Prediction Visualzation. Each row in Fig. 6 shows the input image, ground truth, ZegCLIP’s prediction, and ours. Our method successfully segments unseen categories such as “clouds,” “bushes,” and “playingfield,” which are missed or mislabeled by ZegCLIP. Notably, in both examples, the unseen class “clouds” is correctly identified by our model, demonstrating better generalization to unseen concepts.
@@ 5 Conclusion
In this work, we address the challenge of zero-shot semantic segmentation in query-based frameworks, where conventional Hungarian matching fails to account for unannotated objects. We propose Split Matching (SM), a novel mechanism that identifies and assigns unseen candidates from unannotated regions via CLIP-guided pseudo masks and class embeddings. Furthermore, we introduce an auxiliary Multi-scale Feature Enhancement (MFE) module to enhance the quality of key-value representations in the transformer decoder. Our method is the first to extend hungarian matching to the zero-shot setting and achieves state-of-the-art performance on two zero-shot segmentation datasets.

### v3

@@ 4.1 Comparison with State-of-the-art Methods
Table 1 shows that our method achieves state-of-the-art performance, outperforming existing methods in terms of overall hIoU and uIoU. Specifically, we obtain the highest uIoU on PASCAL VOC (83.1%), which is a significant margin over ZegCLIP (77.8%) and OT-Seg (78.1%), demonstrating stronger generalization to unseen classes. Importantly, our method also achieves the best hIoU on both datasets, indicating more balanced segmentation. Although our sIoU on VOC is slightly lower than transformer-based counterparts, this reflects our model’s ability to mitigate seen-class bias.
@@ 4.2 Ablation Studies
To evaluate the effectiveness of our method, we conduct ablation studies on COCO-Stuff for 40K iterations using ResNet-50 as the backbone, with all hyperparameters unchanged.
@@
Ablations on Unseen Class Embedding. Table 4 presents an ablation study on the design of unseen class embeddings $\mathbf{C}_{u}$. Using raw dense features $\mathbf{F}$ from the backbone provides a reasonable baseline (uIoU: 32.1), while adding an MLP projection degrades performance, likely due to semantic distortion. In contrast, using CLS tokens from pseudo-masked regions yields the best results (uIoU: 36.4), suggesting that region-level CLS tokens better preserve semantic alignment. This also enables direct concatenation with seen-class text embeddings, maintaining cross-modal consistency without modality mismatch.
Ablations on structure of MFE. Table 4 presents the ablation study on the structure of the MFE module by comparing different normalization strategies: No Normalization (No Norm), Batch Normalization (BN), and Group Normalization (GN). Among the three, GN achieves the best performance across all metrics, yielding the highest hIoU, sIoU, and uIoU.
@@ 4.3 Qualitative Analysis
Visualization of candidate queries. Fig. 6 visualizes the predictions of candidate queries. Notably, these queries successfully activate on previously unannotated regions, enabling the model to localize unseen classes such as carrot and cow. This demonstrates that candidate queries can effectively discover latent classes and assign semantically correct class labels, even without explicit supervision. Prediction Visualzation. Each row in Fig. 6 shows the input image, ground truth, ZegCLIP’s prediction, and ours. Our method successfully segments unseen classes such as “clouds”, “bushes”, and “playingfield”, which are missed or mislabeled by ZegCLIP. Notably, in both examples, the unseen class “clouds” is correctly identified by our model, demonstrating better generalization to unseen concepts. Visualization of query distribution. Fig. 8 shows seen (blue), candidate (orange), and random (green) queries in the embedding space. Random queries occupy a concentrated region (red box) that complements areas underrepresented by seen and candidate queries, improving coverage of unseen classes. Fig. 8 appears sparser in the lower-left region, explaining the performance gain from random queries by increasing feature density.
Figure 7: T-SNE visualization among all three types of queries.
Figure 8: T-SNE visualization of seen and candidate queries.
@@ 5 Conclusion
In this paper, we propose Split Matching (SM), a novel decoupled assignment strategy tailored for query-based models in ZSS. By separating queries into seen and candidate groups and optimizing them with respect to annotated and unannotated regions, SM effectively mitigates the seen-class bias caused by incomplete supervision. To further facilitate the discovery of unseen classes, we leverage CLIP-derived pseudo masks and region-level embeddings, and introduce a Multi-scale Feature Enhancement (MFE) module to refine spatial representations. Additionally, we incorporate a Random Query (RQ) strategy during inference to improve query diversity and coverage of unannotated regions. Extensive experiments on standard ZSS benchmarks demonstrate that our approach achieves state-of-the-art results.
Acknowledgment
Support for this work was given by the Toyota Motor Corporation (TMC) and JSPS KAKENHI Grant Number 23K28164 and JST CREST Grant Number JPMJCR22D1. However, note that this paper solely reflects the opinions and conclusions of its authors and not TMC or any other Toyota entity. Computations are done on the supercomputer “Flow” at the Information Technology Center, Nagoya University.